OpenStudy (anonymous):

What is the radius of a circle with a circumference of 14.13 meters? Use 3.14 for pi and show work

OpenStudy (anonymous):

r = C/(2π) r = 14.13/(6.28) = 2.25 meters

OpenStudy (anonymous):

i'm i right? ^

OpenStudy (ash2326):

Yeah you're right:)